As treasurer of the Dragon Corporation, Henry Lee is worried about his bad debt ratio, which is currently running at 7%. The company currently has $40 million in sales. He believes that imposing a more stringent credit policy might reduce sales by 10% and reduce the bad debt ratio to 3%. If the cost of goods sold is 70% of the selling price, should Mr. Lee adopt the more stringent policy?
Answer #1

Existing policy

Currently sales $ 40 Million

Cost of goods sold is 70 % of the selling price

Gross profit = 100-70 = 30%

Profit After considering the Bad debt ratio = 30% - 7 % = 23%,

That is $ 40 * 23% = $ 9.2 Millions

New Policy

If Mr. Lee adopt the more stringent policy

He believes that imposing a more stringent credit policy might reduce sales by 10% and reduce the bad debt ratio to 3%

Sales would be (reduce sales by 10%) = $ 40 * 90% = $ 36 millions

Profit After considering the Bad debt ratio (reduce the bad debt ratio to 3%) = 30% - 3 % = 27%,

That is $ 36 * 27% = $ 9.72 Millions

Its better to Adopt the more stringent policy Because profit will increase by = $ 0.52 Millions. ($ 9.72 -$ 9.2 )